# On the Use of WGANs for Super Resolution in Dark-Matter Simulations

### v1

@@ Abstract
Super-resolution techniques have the potential to reduce the computational cost of cosmological and astrophysical simulations. This can be achieved by enabling traditional simulation methods to run at lower resolution and then efficiently computing high-resolution data corresponding to the simulated low-resolution data. In this work, we investigate the application of a Wasserstein Generative Adversarial Network (WGAN) to increase the particle resolution of dark-matter-only simulations, reproducing and building on prior results. Our WGAN models successfully generate high-resolution data with summary statistics, including the power spectrum and halo mass function, that closely match those of true high-resolution simulations. We also identify a limitation of the WGAN model in the form of smeared features in the generated high-resolution snapshots, particularly in the shapes of dark-matter halos.
@@
Machine Learning, Super Resolution, Cosmological Simulations, Dark-Matter Simulations
@@ 1 Introduction
Super resolution (SR) refers to the task of artificially increasing the resolution of data. The computer science community has worked on this problem for several decades and, as a result, developed several families of techniques for computing high-quality solutions to the problem (e.g. Lepcha et al., 2023). Although the initial use case for most of these techniques was image enhancement, all of them can, in principle, be used to improve the resolution of physics simulations (e.g. Fukami, Fukagata & Taira, 2019; Xie et al., 2017; Byeong-Sun et al., 2021). Instead of enhancing the red, green and blue (RGB) channels of a two-dimensional image, one can think about enhancing various field channels (e.g. a density channel, temperature channel, velocity component channels etc.) of a three-dimensional simulation snapshot. Of particular interest to the authors are applications of these techniques to improving the resolution of dark-matter cosmological simulations.
@@
The main reason to consider SR techniques for astrophysics simulations is their potential to significantly reduce the computational cost of producing high-resolution (HR) data. While HR datasets are crucial for theoretical work in many areas of astrophysics, the simulations needed to produce them are computationally expensive. More broadly, fast approximation techniques for structure formation simulations are essential for observational cosmology surveys. These surveys require numerous realizations of cosmic large-scale structures on Gigaparsec scales, while also capturing the distribution of galaxies and their satellites on scales ranging from tens to hundreds of kiloparsecs (e.g. Pujol et al., 2017; Ding et al., 2022). Thus, a suitable balance of resolution regularly needs to be found in order to produce meaningful results while avoiding impractical or intractable simulations - a task which is not always feasible. SR methods offer an approach to bridge the gap between low-resolution (LR) and HR simulations, allowing researchers to run cheaper LR simulations and recover HR data from them. For instance, while not quite SR, similar methods based on deep learning allowed Jacobus et al. (2024) to make HR corrections to LR simulations without increasing its resolution. This allowed them to produce high-quality data representing a gigaparsec hydrodynamic volume, far beyond the capabilities of traditional simulation methods running on current supercomputers.
@@
In a trilogy of papers, (Li et al., 2021; Ni et al., 2021; Zhang et al., 2024) demonstrated the impressive capabilities of a generative adversarial neural network (GAN) (Goodfellow et al., 2014) in enhancing the particle resolution of dark-matter-only simulations. This was done by first converting particle position data into a displacement field with respect to an initial grid. A GAN model, inspired by a popular design known as styleGAN (Karras, Samuli & Timo, 2019; Karras et al., 2020) for image generation, was then used to enhance the resolution of this displacement field. The enhanced displacement field is then converted back to particle position data with a higher particle count. The models they developed allowed them to increase the particle resolution of their simulation snapshots by a factor of 512 over a range of redshifts. This increase in resolution enabled them to model HR simulations well enough to reproduce power spectra, halo mass functions and halo merger trees.
In this work, we reproduce some early results in this area for smaller enhancement factors and report on a limitation of the proposed GAN model in predicting HR data from corresponding LR data. Namely, we show that the model has difficulty in reproducing the shape of dark-matter halos and show how the discrepancy with HR simulations can be highlighted. We also present results from attempts to address this limitation.
@@
The paper is structured as follows: in section 2 we introduce our methods along with the datasets used and section 3 lists our main results. We summarise and conclude in section 4. Throughout we use a WMAP9 cosmology (Hinshaw et al., 2013): $(\Omega_{cdm},\Omega_{\Lambda},\Omega_{b},h)=(0.276,0.724,0,0.703)$.
@@ 2.1 Dataset Details
The training and validation datasets used in this work were created from a set of dark-matter-only simulation pairs, all of comoving volume $(100\leavevmode\nobreak\ {h}^{-1}\mbox{Mpc})^{3}$. Each simulation pair consists of a LR and a HR simulation of $64^{3}$ and $128^{3}$ particles respectively, with corresponding particle masses of $4.2\times 10^{11}M_{\odot}$ and $5.2\times 10^{10}M_{\odot}$. Initial conditions were generated using the MUSIC initial conditions generator (Hahn & Abel, 2011). For each simulation pair, we fixed the random seed used to generate their initial conditions. Each pair is therefore a low- and high-resolution realization of the same cosmological volume. All our simulations were run from redshift $z=99$ down to $z=0$ with a gravitational softening length of 1/30th the mean inter-particle separation using the SWIFT simulation code (Schaller et al., 2024). Snapshots were taken at redshift 0 and a set of 16 simulation pairs was used to create the training dataset while a single simulation pair was used for the validation dataset.
Both datasets consisted of pairs of LR and HR patches of dark-matter displacement fields. These were created by first converting the position data, of dark-matter particles in each snapshot, into displacement fields. This was achieved by computing the relative position of each particle with respect to the points of the regular grid that define their initial position (at $z=99$). Displacement fields were stored as multi-dimensional arrays, or tensors, with shape $(3,N,N,N)$ where the first dimension corresponds to the $x$, $y$ and $z$ components, or channels, of the displacement field. The next three dimensions correspond to the $x$, $y$ and $z$ indices of points on the $N\times N\times N$ regular grid. Here, $N$ is 64 and 128 for the LR and HR patches respectively. These arrays were then segmented into patches, yielding a set of 64 LR patches of shape $3\times 20^{3}$, and $3\times 32^{3}$ HR counterparts, for each snapshot. LR patches are taken as $3\times 16^{3}$ patches with 2 cells of padding on each side. Correspondingly, the $3\times 32^{3}$ HR patches should be taken as HR copies of the inner $3\times 16^{3}$ regions of the LR patches. Both the HR patches and the inner $16^{3}$ regions of the LR patches represent cubes with a spatial length of $25\ \rm{h^{-1}}\mbox{Mpc}$ (comoving). Random permutations of the $\rm{xyz}$-axes were also used as a data augmentation technique to artificially expand the dataset size during training.
@@ 2.2 Neural Network Details
Following the work of Li et al. (2021), a Wasserstein generative adversarial neural network model (WGAN) (Arjovsky, Chintala & Bottou, 2017) was used to enhance dark-matter data. WGANs consist of two neural network models: a generator model and a critic, or discriminator, model. The generator model is responsible for producing synthetic HR versions of any LR data passed to it, while the critic model is used to assign a score to any HR data it is given. The score assigned by the critic model should be taken as a measure of how much the data resembles real HR simulation data. Below we outline the neural network architectures used for both of these models.
@@ 2.2.1 Generator Model
Figure 1: In the above graphics, tensors are denoted by gray boxes with the tensor shape shown inside. All tensors are rank 4 and have a shape $C\times N^{3}$ where $C$ is the number of channels of the data and $N$ is the number of cells in each of the $\rm{xyz}$ directions. Operations acting on tensor data are denoted by rounded boxes with the name of the operation inside. Key parameters are also shown in the operation boxes: the kernel size for convolutional layers, the crop size for crop operations and the scale factor for trilinear interpolation operations. Where convolutional layers also change the number of channels, the new number of channels is also given in the corresponding box. On the left is a graphical representation of the generator architecture. The input tensor $X$, containing LR data, is passed to both a primary and an auxiliary data stream, indicated by blue and red arrows respectively. Data in both streams are processed and upscaled by a sequence of H-blocks. The auxiliary stream is also processed by an initial convolutional block to expand the number of channels in the stream. The output tensor of the primary stream is cropped before it is taken as the output $Y$ of the entire model. Each H-block additionally takes two tensors, containing Gaussian noise, as input. The collection of all such tensors is denoted by $Z$ and referred to as a latent variable. The graphic on the right shows the architecture of a H-block layer.
In addition to the two main inputs, each H-block also takes standard Guassian noise as extra input. More precisely, when a H-block receives a primary input tensor with a shape $(3,N,N,N)$, two additional tensors are provided as input: one with shape $(1,N,N,N)$ and another with shape $(1,2N-2,2N-2,2N-2)$. Both of these extra tensors contain samples drawn from a standard normal distribution. This noise is used by the H-block to synthesize new HR details to be added to the upscaled LR data. In keeping with the terminology used in the literature on GANs, the collection of all such noise tensors passed to the H-blocks is referred to as a sample $z$ from a latent space $Z$ for the generator. Denoting the generator by $G$ and the LR data to be upscaled by $X$, then for a given sample $z$ from the latent space, the enhanced data can be written as by $Y=G(X,z)$. Furthermore, for two distinct samples from the latent space $z_{1}$ and $z_{2}$, the generator outputs $Y_{1}=G(X,z_{1})$ and $Y_{2}=G(X,z_{2})$ can be taken as distinct HR realizations of the same LR input $X$.
@@
To outline the architecture of the H-block layer (depicted in Fig. 1), we note it is responsible for two key tasks. Firstly, it increases the resolution of data in both the primary and auxiliary streams. Data tensors with spatial size $N$ are upscaled, by a single H-block, to a size $42N-$ 4. This is done in the axillary stream using a trilinear interpolation operation to increase the spatial dimensions by a factor of 2, followed by a pair of convolutional layers which remove 1 cell of padding each from the edges of the tensor. In the primary data stream, a trilinear interpolation operation is again used to increase the spatial dimensions by a factor of 2 and then a crop operation removes 2 cells of padding from the edges of the tensor to keep the spatial dimensions in both streams aligned. Secondly, the H-block injects noise into the feature channels of the auxiliary stream. This is done by concatenating noise with the data in the auxiliary stream at two points: before the interpolation operation and before the second convolutional layer as depicted in Fig. 1. The noise added to the data is the same noise passed to the H-block as input as discussed above. Finally, a projection operation is performed by a convolutional layer that maps the auxiliary data into a 3 channel data tensor to be added to the primary data before being returned as output. We can think of this operation as adding any HR details synthesized by the auxiliary stream to the primary output.
Each H-block both increases the spatial dimensions of the input data given to its’ primary and auxiliary streams and halves the number of channels used by the auxiliary stream. More precisely, for auxiliary and primary inputs with shapes $(C,N,N,N)$ and $(3,N,N,N)$ respectively, the auxiliary and primary outputs of the H-block have shapes $(C/2,2N-4,2N-4,2N-4)$ and $(3,2N-4,2N-4,2N-4)$ respectively. The number of H-blocks used in the generator model thus determines the factor by which LR data is enhanced. As mentioned, in this work we focus on enhancing LR data by a factor of 2. We therefore use a single H-block in our generator model and pad our LR data with 2 cells on each side to account for the boundary cells consumed by the convolutional layers in the H-block. The spatial dimensions of the generator output were cropped on each side by 2 cells.
@@ 2.3 Training Details
Once the critic step completes, the generator step is performed. During the generator step, more latent variables $z$ are sampled from the latent space and the generator is used again to create fake HR counterparts to the LR data $x$ in the batch. This data is then prepared as input for the critic model, $X_{G(x,z)}$, by similarly pairing it with upscaled LR data and associated density fields as in the critic step. The following loss value is then computed, the gradient of which is used by the generator optimizer to update the generator parameters:
@@
After this update step, we expect intuitively that the generator improves at generating data that receive a high score from the critic model and thus better resembles real HR data.
@@ 3 Results
We now highlight results from enhancing the resolution of a snapshot from a dark-matter-only simulation. The snapshot was enhanced using WGANs that were trained as described in the previous section. Note that, while all models used in this work were trained on the training set discussed in 2.1, the snapshot that was enhanced here was taken from the validation set. That is, the simulation data used to produce the plots that follow were not used to train any of the models. The following results therefore demonstrate the ability of the trained models to generalize and enhance data they have not seen during training. We also note that all models used in this work were trained on a NVIDIA H100 GPU with 64GB of memory.

### v2

# On the Use of WGANs for Super–Resolution in Dark-Matter Simulations

@@ Abstract
Super-resolution techniques have the potential to reduce the computational cost of cosmological and astrophysical simulations. This can be achieved by enabling traditional simulation methods to run at lower resolution and then efficiently computing high-resolution data corresponding to the simulated low-resolution data. In this work, we investigate the application of a Wasserstein Generative Adversarial Network (WGAN) model, previously proposed in the literature, to increase the particle resolution of dark-matter-only simulations. We reproduce prior results, showing the WGAN model successfully generates high-resolution data with summary statistics, including the power spectrum and halo mass function, that closely match those of true high-resolution simulations. However, we also identify a limitation of the WGAN model in the form of smeared features in generated high-resolution data, particularly in the shapes of dark-matter halos and filaments. This limitation points to a potential weakness of the proposed WGAN-based super-resolution method in capturing the detailed structure of halos, and underscores the need for further development in applying such models to cosmological data.
@@
Machine Learning, Super-Resolution, Cosmological Simulations, Dark-Matter Simulations
@@ 1 Introduction
Super-resolution (SR) refers to the task of artificially increasing the resolution of data. The computer science community has worked on this problem for several decades and, as a result, developed several families of techniques for computing high-quality solutions to the problem (e.g., Lepcha et al., 2023). Although the initial use case for most of these techniques was image enhancement, all of them can, in principle, be used to improve the resolution of physics simulations (e.g., Fukami, Fukagata & Taira, 2019; Xie et al., 2017; Byeong-Sun et al., 2021). Instead of enhancing the red, green and blue (RGB) channels of a two-dimensional image, one can think about enhancing various field channels (e.g., a density, temperature, velocity component etc.) of a three-dimensional simulation snapshot. Of particular interest to the authors are applications of these techniques to improving the resolution of dark-matter cosmological simulations.
@@
The main reason to consider SR techniques for astrophysics simulations is their potential to significantly reduce the computational cost of producing high-resolution (HR) data. While HR datasets are crucial for theoretical work in many areas of astrophysics, the simulations needed to produce them are computationally expensive. More broadly, fast approximation techniques for structure formation simulations are essential for observational cosmology surveys. These surveys require numerous realizations of cosmic large-scale structures on Gigaparsec scales, while also capturing the distribution of galaxies and their satellites on scales ranging from tens to hundreds of kiloparsecs (e.g., Pujol et al., 2017; Ding et al., 2022). Thus, a suitable balance between resolution and computational cost regularly needs to be found in order to produce meaningful results while avoiding impractical or intractable simulations – a task which is not always feasible. SR methods offer a potential approach to bridge the gap between low-resolution (LR) and HR simulations, allowing researchers to run cheaper LR simulations and recover HR data from them. For instance, while not quite SR, similar methods described in Jacobus et al. (2024) allowed the authors to make HR corrections to LR simulations without increasing its resolution. This enabled them to produce high-quality data representing a gigaparsec hydrodynamic volume, far beyond the capabilities of traditional simulation methods running on current supercomputers.
@@
In a trilogy of papers, (Li et al., 2021; Ni et al., 2021; Zhang et al., 2024) proposed a method based on a generative adversarial neural network (GAN) (Goodfellow et al., 2014) for enhancing the particle resolution of dark-matter-only simulations. Their approach involved converting particle position data into a displacement field with respect to an initial grid. A Wasserstein GAN model (WGAN) (Arjovsky, Chintala & Bottou, 2017), inspired by a popular design known as styleGAN (Karras, Samuli & Timo, 2019; Karras et al., 2020) for image generation, was then used to enhance the resolution of this displacement field. The enhanced displacement field is then converted back to particle position data with a higher particle count. Their models achieved up to a 512-fold increase in particle resolution across a range of redshifts and reproduced several key statistical properties of high-resolution simulations, including power spectra, halo mass functions, and halo merger trees.
In this work, we reproduce earlier results for modest super-resolution enhancement factors and highlight a key limitation of the WGAN-based approach in predicting high-resolution data from corresponding low-resolution inputs. Specifically, we find that the model struggles to accurately reproduce the shapes of dark-matter halos and filaments, leading to noticeable discrepancies when compared with true high-resolution simulations. We further investigate this issue and present initial results from efforts to mitigate the problem.
@@
The paper is structured as follows: in section 2 we introduce the methods along with the datasets used, and section 3 lists our main results. We summarise and conclude in section 4. Throughout we use a WMAP9 cosmology (Hinshaw et al., 2013): $(\Omega_{m},\Omega_{\Lambda},\Omega_{b},h)=(0.276,0.724,0,0.703)$.
@@ 2.1 Dataset Details
The training, validation and test datasets used in this work were created from a set of dark-matter-only simulation pairs, all of comoving volume $(100\leavevmode\nobreak\ {h}^{-1}\mbox{Mpc})^{3}$. Each simulation pair consists of an LR and an HR simulation of $64^{3}$ and $128^{3}$ particles respectively, with corresponding particle masses of $4.2\times 10^{11}M_{\odot}$ and $5.2\times 10^{10}M_{\odot}$. Initial conditions were generated using the MUSIC2–monofonIC111bitbucket.org/ohahn/monofonic/src/master/ initial conditions generator (Michaux et al., 2020; Hahn, Rampf & Uhlemann, 2021). For each simulation pair, we fixed the random seed used to generate their initial conditions. Each pair therefore represents both a low- and a high-resolution realization of the same cosmological volume. All our simulations were run from redshift $z=99$ down to $z=0$ with a gravitational softening length of 1/30th the mean inter-particle separation using the SWIFT simulation code (Schaller et al., 2024). Snapshots were taken at redshift 0 and a set of 16 simulation pairs was used to create the training dataset while a single simulation pair each was used for both validation and test datasets.
All datasets consisted of pairs of LR and HR patches of dark-matter displacement fields. These were created by first converting the position data, of dark-matter particles in each snapshot, into displacement fields. This was achieved by computing each particle’s displacement from its initial grid position. Displacement fields were stored as multi-dimensional arrays, or tensors, with a shape $(3,N,N,N)$ where the first dimension corresponds to the $x$, $y$ and $z$ components, or channels, of the displacement field. The next three dimensions correspond to the $x$, $y$ and $z$ indices of points on the $N\times N\times N$ regular grid. Here, $N$ is 64 and 128 for the LR and HR patches respectively. These arrays were then segmented into patches, yielding a set of 64 LR patches of shape $3\times 20^{3}$, and $3\times 32^{3}$ HR counterparts, for each snapshot. This resulted in $64\times 16=1024$ LR-HR patch pairs in the training set, and 64 patch pairs in both the validation and test sets. We take the LR patches to be $3\times 16^{3}$ patches with 2 cells of padding on each side, making the total patch size $3\times 20^{3}$. Correspondingly, the $3\times 32^{3}$ HR patches should be taken as HR copies of the inner $3\times 16^{3}$ regions of the LR patches. Both the HR patches and the inner $16^{3}$ regions of the LR patches represent cubes with a spatial length of $25\ \rm{h^{-1}}\mbox{Mpc}$ (comoving). Random permutations of the $\rm{xyz}$-axes were also used as a data augmentation technique to artificially expand the dataset size during training.
@@ 2.2 Neural Network Details
Following the approach of Li et al. (2021), a Wasserstein generative adversarial neural network model with a gradient penalty (WGAN) (Arjovsky, Chintala & Bottou, 2017; Gulrajani et al., 2017) was used to enhance dark-matter data. WGANs are widely regarded as a more stable and robust alternative to classical GAN models, mitigating common issues such as mode collapse and training instability. They consist of two neural network models: a generator model and a critic, or discriminator, model. The generator model is responsible for producing synthetic HR versions of any LR data passed to it, while the critic model is used to assign a score to any HR data it is given. The score assigned by the critic model should be taken as a measure of how much the data resembles real HR simulation data. Below we outline the neural network architectures we used for both of these models.
@@ 2.2.1 Generator Model
Figure 1: The graphics above illustrate the tensor operations used by the genertaor model. Tensors are denoted by gray boxes with thick borders, with their shape indicated inside. All tensors are rank 4 and have a shape $C\times N^{3}$ where $C$ is the number of data channels and $N$ is the number of cells in each of the $\rm{xyz}$ directions. Operations on tensors are shown as rounded boxes with the name of the operation inside. Key parameters are also shown in the these boxes: the kernel size for convolutional layers, the crop size for crop operations and the scale factor for trilinear interpolation operations. If a convolutional layer changes the number of channels, the new number of channels is also given. The generator architecture is depicted on the left. The input tensor $X$ (LR data), is passed to both a primary and an auxiliary data stream, indicated by blue and red arrows respectively. Data in both streams are processed and upscaled by a sequence of H-blocks. The auxiliary stream is also processed by an initial convolutional block to expand the number of channels in the stream. The output of the primary stream is cropped and taken as the output $Y$ of the model. Each H-block additionally takes two tensors, containing Gaussian noise, as input. The collection of all such tensors is denoted by $Z$ and constitutes the latent variable for the model. The graphic on the right shows the internal structure of a H-block.
In addition to the two main inputs, each H-block also takes standard Gaussian noise as extra input. More precisely, when a H-block receives a primary input tensor with a shape $(3,N,N,N)$, two additional tensors are provided as input: one with shape $(1,N,N,N)$ and another with shape $(1,2N-2,2N-2,2N-2)$. Both of these extra tensors contain samples drawn from a standard normal distribution. This noise is used by the H-block to synthesize new HR details to be added to the upscaled LR data. In keeping with the terminology used in the literature on GANs, the collection of all such noise tensors passed to the H-blocks is referred to as a sample $z$ from a latent space $Z$ for the generator. Denoting the generator by $G$ and the LR data to be upscaled by $X$, then for a given sample $z$ from the latent space, the enhanced data can be written as $Y=G(X,z)$. Furthermore, for two different samples from the latent space $z_{1}$ and $z_{2}$, the generator outputs $Y_{1}=G(X,z_{1})$ and $Y_{2}=G(X,z_{2})$ can be taken as non-identical HR realizations of the same LR input $X$. These HR realizations differ in the small-scale details, particularly in regions where the LR data underdetermines the HR structure, while remaining consistent with the large-scale structure set by LR input.
@@
To outline the architecture of the H-block layer (depicted in Fig. 1), we note it is responsible for two key tasks. Firstly, it increases the resolution of data in both the primary and auxiliary streams. Data tensors with spatial size $N$ are upscaled, by a single H-block, to a size $42N-$ 4. This is done in the auxiliary stream using a trilinear interpolation operation (linear interpolation in three dimensions) to increase the spatial dimensions by a factor of 2, followed by a pair of convolutional layers which remove 1 cell of padding each from the edges of the tensor. In the primary data stream, a trilinear interpolation operation is again used to increase the spatial dimensions by a factor of 2 and then a crop operation removes 2 cells of padding from the edges of the tensor to keep the spatial dimensions in both streams aligned. Secondly, the H-block injects noise into the feature channels of the auxiliary stream. This is done by concatenating noise with the data in the auxiliary stream at two points: before the interpolation operation and before the second convolutional layer as depicted in Fig. 1. The noise added to the data is the same noise passed to the H-block as input as discussed above. Finally, a convolutional layer performs a projection operation that maps the auxiliary data into a 3-channel tensor. This tensor is then added to the primary data before being returned as output. We can think of this operation as adding any HR details synthesized by the auxiliary stream to the primary output.
Each H-block both increases the spatial dimensions of the input data given to its primary and auxiliary streams and halves the number of channels used by the auxiliary stream. More precisely, for auxiliary and primary inputs with shapes $(C,N,N,N)$ and $(3,N,N,N)$ respectively, the auxiliary and primary outputs of the H-block have shapes $(C/2,2N-4,2N-4,2N-4)$ and $(3,2N-4,2N-4,2N-4)$ respectively. The number of H-blocks used in the generator model thus determines the factor by which LR data is enhanced. As mentioned, in this work we focus on enhancing LR data by a factor of 2. We therefore use a single H-block in our generator model and pad our LR data with 2 cells on each side to account for the boundary cells consumed by the convolutional layers in the H-block. The spatial dimensions of the generator output were cropped on each side by 2 cells.
@@ 2.3 Training Details
Once the critic step completes, the generator step is performed. During the generator step, more latent variables $z$ are sampled from the latent space and the generator is used again to create fake HR counterparts to the LR data $x$ in the batch. This data is then prepared as input for the critic model, $X_{G(x,z)}$, by similarly pairing it with upscaled LR data and associated density fields as in the critic step. The following loss value is then computed, and its gradient is used by the generator optimizer to update the generator parameters:
@@
After this update step, the generator is expected to improve in generating data that receives a higher score from the critic model, and thus better resembles real HR data, as judged by the critic.
@@ 3 Results
We now present results from enhancing the resolution of a snapshot from a dark-matter-only simulation. We first show that the model can produce data with expected summary statistics and then present a limitation of the approach: its failure to reproduce the morphology of dark-matter halos and filaments. Note that, while all models used in this work were trained on the training set discussed in 2.1, the snapshot that was enhanced here was taken from the test set. That is, the simulation data used to produce the plots that follow were not used to train any of the models. We also note that all models used in this work were trained on a NVIDIA H100 GPU with 64GB of memory.